OpenStudy (anonymous):

Solve for the variable k/3=16/24 a. 1/2 b. 2/3 c. 2 d. 6

OpenStudy (anonymous):

k/3=2/3 k=2

OpenStudy (anonymous):

\[k/3=16/24\] multiply both sides by three to isolate k

OpenStudy (paki):

k/3=16/24 k= (16/24)*3 k=48/24 k=2